OpenStudy (fibonaccichick666):

@agent0smith The 3 main kinetic equations in basic physics. Do they all derive from each other?

OpenStudy (raffle_snaffle):

Yes

OpenStudy (raffle_snaffle):

You mean kinematics?

OpenStudy (agent0smith):

They're derived with calculus starting with \( a(t) = a\) (for constant acceleration). The v^2 = vo^2 + 2a(x-xo) comes from combining a couple of them.

OpenStudy (raffle_snaffle):

If you want I can show you Fibchic.

OpenStudy (fibonaccichick666):

Could you @raffle_snaffle ? I would like to see it. Or just give me a start. I can go from the basic start probably. I thought it was something like \[v_{avg}=\frac{v_f-v_0}{2}\] then doing something to get \[v_f^2=v_0^2+2ax\]

OpenStudy (agent0smith):

Integrating \(a(t) = a\) twice gives the below equations. To get \(\large v_f^2=v_o^2+2a(x_f-x_o)\)... \(\Large v_f=v_o + at\) <=== solve for t \(\Large x_f =x_o+ v_o t + \frac{ 1 }{ 2 }at^2 \) <=== substitute it for t in here Simplify and you'll have that equation.

OpenStudy (fibonaccichick666):

ah, I meant isn't there another way other than the integration. I remember my physics teacher doing some sort of algebraic manipulation. The calc version isn't bad

OpenStudy (agent0smith):

To get what exactly?